OpenStudy (ragingsquirrel):

A wall is 12 feet long and 8 feet tall. There is a square window 4 feet long in the wall. What is the area of the wall surface? 96 ft2 92 ft2 80 ft2 32 ft2

OpenStudy (chriss):

12 * 8=96 4 * 4=16 96-16=80 80 square feet

OpenStudy (anonymous):

(12x8 wall)-(4x4 window) = 80 sqft.